OpenStudy (angelinaxoxo1):

Write 15 to 40 as a ratio in simplest form.

OpenStudy (gavinsauntie):

http://www.virtualnerd.com/middle-math/ratios-proportions-percent/ratios-rates/ratio-simplest-form-example Hope this helps! It always helped me when I would put ratios in simplest form.

OpenStudy (michele_laino):

we have these factorizations: \[\begin{gathered} 15 = 3 \times 5 \hfill \\ 40 = {2^3} \times 5 \hfill \\ \end{gathered} \] so, what is the great common divisor?

OpenStudy (michele_laino):

we have: greatest common divisor=5 so, please divide both numerator and denominator by 5

OpenStudy (michele_laino):

15:5=...? 40:5=...? please complete